OpenStudy (anonymous):

which of the following is equivalent to x^5y^2/xy^2 when y is not equal to 0 and y is not equal to 0 x^6y^5 x^5y x^4y x^4

OpenStudy (superdavesuper):

u can simplify by using 2 eqns: x^a / x^b = x^(a-b) n x^0=1 simplify both x^ and y^ to get the ans. :)

OpenStudy (anonymous):

I think the answer is x ^4

OpenStudy (superdavesuper):

bingo @sk8ter02468 :)

OpenStudy (anonymous):

Thank you

OpenStudy (superdavesuper):

welcome